### v1

# Distilling Event Sequence Knowledge From Large Language Models

Abstract
Event sequence models have been found to be highly effective in the analysis and prediction of events. Building such models requires availability of abundant high-quality event sequence data. In certain applications, however, clean structured event sequences are not available, and automated sequence extraction results in data that is too noisy and incomplete. In this work, we explore the use of Large Language Models (LLMs) to generate event sequences that can effectively be used for probabilistic event model construction. This can be viewed as a mechanism of distilling event sequence knowledge from LLMs. Our approach relies on a Knowledge Graph (KG) of event concepts with partial causal relations to guide the generative language model for causal event sequence generation. We show that our approach can generate high-quality event sequences, filling a knowledge gap in the input KG. Furthermore, we explore how the generated sequences can be leveraged to discover useful and more complex structured knowledge from pattern mining and probabilistic event models. We release our sequence generation code and evaluation framework, as well as corpus of event sequence data.
1 Introduction
Building probabilistic models from event sequence data has numerous practical applications across different domains when plentiful high-quality event data is available. For example, in Finance, event models can be used to predict stock market trends and make informed investment decisions. In Healthcare, event models can help identify patterns and correlations in patient data to improve diagnoses and treatment plans. In the field of Cybersecurity, these models can be used to detect and prevent potential cyber attacks by analyzing the sequence of events leading up to a breach. A common characteristic of the data in these domains is that sequences are clearly associated with an entity (e.g., a company, a person, or a device). There are however other domains where such a clean association between events and entities may not be possible. One such application is news event analysis Radinsky et al. (2012); Cekinel and Karagoz (2022); Du et al. (2022); Hassanzadeh et al. (2022). While various news sources record and describe newsworthy events, it is often not possible to automatically put together coherent sequences of events, because each event may involve multiple topics and actors, and many correlated and unrelated events may be occurring simultaneously or in close proximity
Prior work has addressed this challenge by devising automated mechanisms for extracting narratives Norambuena et al. (2023); Santana et al. (2023), topic detection and tracking Allan (2012), and timeline summarization Gholipour Ghalandari and Ifrim (2020). While these different categories of solutions have been successful in a range of applications, the outcome is inherently noisy and not in the form of structured event sequences useful for the construction of event models. Prior work has shown little success in automatically turning narratives into structured sequences. In this paper, we explore a novel mechanism for creating structured event sequences in such domains, using generative language models.
Figure 1: Examples of frequent patterns mined from LLM-generated event sequence collections with their support (top); and some results from a simple probabilistic event model learned over the sequences that discovers influencing sets of event concepts and estimates probabilities based on historical occurrences (bottom).
Large Language Models Brown et al. (2020); Raffel et al. (2020); Wei et al. (2022) have recently become the dominant paradigm in a range of natural language processing (NLP) tasks Chung et al. (2022) and often beat traditional approaches on a number of challenging tasks, including complex arithmetic reasoning Imani et al. (2023) and open-domain question answering Kamalloo et al. (2023). In this paper, our goal is to examine the capability of LLMs to generate structured event sequences useful for event analysis. Our aim is to use LLMs directly for generation of event sequences, as opposed to improving prior extraction methods that would rely on high-quality and comprehensive sources describing detailed event sequences for a wide variety of events. Our hypothesis is that LLMs trained on large corpora have already gathered the required knowledge of plausible event sequences and therefore can be suitably guided to produce diverse and high-quality event sequences. To effectively distill this knowledge, we use event-related concepts in Wikidata Vrandečić and Krötzsch (2014), a comprehensive general-domain knowledge graph, to guide the sequence generation. This can be viewed as a novel mechanism for knowledge-guided text generation Yu et al. (2022). We then use these generated sequences for pattern mining and learning probabilistic event models, as a way of structuring the underlying knowledge. Figure 1 shows examples from our experiments of patterns mined from an LLM-generated event sequence collection, as a well as a simple model learned from the collection.
In summary, we make the following contributions:
  1. 1.
    We devise a new iterative in-context prompting strategy for generating high-quality event sequences using LLMs. To the best of our knowledge, we are the first to use LLMs to generate structured event sequences for the purpose of analyzing various event models.
  2. 2.
    We compile high-quality event sequences using our generation mechanism, based on a curated set of high-level event concepts (classes) from Wikidata that represent newsworthy events.
  3. 3.
    We develop an evaluation framework and conduct experiments to show the value of LLM-powered event sequence generation on replicating and augmenting knowledge in structured representations such as knowledge graphs.
  4. 4.
    We further demonstrate the practical usefulness of our approach by leveraging downstream pattern mining and probabilistic event models.
Upon publication, we will make our code and generated sequence data publicly available (as a separate dataset) for future research.
2 Methods
We split the task of distilling event sequence knowledge from LLMs into two sub-tasks: (i) generation of new and unique event sequences given an event trigger type; and (ii) mining high-support sequence patterns not already derivable from the knowledge graph, and learning probabilistic models to discover complex event sequence rules. Figure 2 describes our overall framework for carrying out these two objectives. We frame these two subtasks as instances of casual event prediction and event sequence mining, respectively.
Figure 2: An overview of our event sequence generation framework and an example portraying a potential use case. We show that by (1) starting with a sparse knowledge graph such as Wikidata, we can generate targeted event sequences. Owing to the inherent sparsity in the underlying KG, we can (2) use LLMs to carry out a portion of the evaluation (i.e. precision) to select an optimal model. On this new generated sequence dataset, we then (3) apply classical pattern-mining algorithms (i.e. GSP) to identify potentially interesting has_normal-_\__cause and has_normal-_\__effect event sequence chains, and (4) learn summary Markov models (SuMMs) to identify potential influencing events (i.e. parent event types) for particular event types of interest; these are both illustrations of extracting complex structured knowledge from the generated sequences.
In real-world knowledge graphs like Wikidata, such events are represented by a text-based identifier. For example, each event type in Figure 2 corresponds to a high-level newsworthy concept in Wikidata. More importantly, these event concepts are not associated with any relevant time-stamps but often do have causal relations (e.g. has_cause or has_effect) – either directly (earthquake causes tsunami) or through their instances (an earthquake event caused a tsunami event).
2.1 LLMs as Event Sequence Generators
We first analyze the ability of LLMs to model event prediction as a conditional generation task under zero and iterative few-shot settings. Our targets are linearized sequences of event concepts. We begin by prompting a large language model (with in-context exemplars) with an event trigger $y_{1}$ to generate the next concept from a defined set of labels, and repeat this process until we get a sequence of desired length. Formally, given an event trigger $\mathcal{T}$, we model the probability of generating linearized string $y$ of length $T$ containing $N$ unique event concepts that follow $\mathcal{T}$ in sequence:
$p_{\text{LM}}(y|\mathcal{T})=\prod_{t=2}^{T}p(y_{t}|\mathcal{T},y_{<t-1})$
This is the standard conditional language modeling objective. We try multiple prompting techniques and qualitatively observe optimal results with an iterative in-context few-shot prompting strategy (Figure 3). Specifically, we start with a set of six randomly selected examples of the form – “What usually follows event $X$?”. Based on the model output ($Y$) from a pre-specified vocabulary, we append this same example to the original prompt in conjunction (i.e. “What usually follows event $X$ and $Y$?”) with ICL examples of the same form. As shown in Figure 3, our iterative technique serves dual purposes: (i) it leverages in-context learning; and (ii) eliminates the need for implementing complex resolvers to post-process model outputs. We repeat this process until a sequence of a desired minimum length $m$ is achieved (in our experiments, $m=3$) or we’ve exhausted a maximum number of tries ($k=10$ in our experiments) to generate an in-domain event type.
Figure 3: Illustration of our approach to elicit has_normal-_\__effect and has_normal-_\__cause events given a label of interest. Use of instructional in-context exemplars substantially reduces the need for post-processing LLM output in addition to constraining the output label space.
2.2 Event Concepts
We curate a new dataset of high-level event concepts (classes) from Wikidata that represent newsworthy events. To do so, we query Wikidata for event concepts that have links to Wikinews articles and are instances of classes that are a subclass of the occurrence class, i.e., indicating they are newsworthy event classes. We gathered $50$ top-level classes for these event concepts, each having multiple labels (e.g. conflict $\rightarrow$ conflict (psychological), dispute, disagreement, etc.). This yielded a total of $202$ unique event labels for the $50$ top-level classes. Most of these event concepts have some causal relations (i.e. has_cause or has_effect). We use these relation pairs as in-context exemplars to create our prompts. Full length prompts used in all our experiments are provided in Appendix LABEL:sec_apx:event_seq_prompts.
Algorithm 1 Generalized Sequential Pattern Algorithm
Obtain sequences $(<x>)$ as $len-1$ candidates
$F_{1}=$ set of $len-1$ sequential patterns
$k=1$
[…]
     Generate $C_{k+1}$triangleright $len-(k+1)$ candidates from $F_{k}$
     if $C_{k+1}\neq\emptyset$  then
         Scan all sequences
         Identify $F_{k+1}$triangleright $len-(k+1)$ sequence patterns
     end if
     $k=k+1$
end while
2.3 Mining New Patterns
In order to derive new, unseen relationships (has_cause or has_effect) between the extracted event classes, we use a large pretrained language model to generate new event sequences and then use classical frequent itemset mining algorithms like GSP Srikant and Agrawal (1996) and SPADE Zaki (2004) to derive new high support patterns. This aspect of the overall workflow is illustrated in the top right section of Figure 2, where rules are mined from event sequences.
To generate new event sequences in a zero-shot setting, we start with an event-trigger (e.g. a concept like workplace accident), create a prompt with instructions describing desired relationships and a few in-context exemplars (ICL prompt), and constrained the model output to the original $50$ event concept labels to generate the next event in the sequence. (The full text of the prompt is provided in the supplementary material.) We append the model output to another ICL prompt with conjunctive event examples (i.e. questions of the form “What happens after X, Y, and Z”). We repeat this process until we reach a pre-defined maximum sequence length (in this case, $10$) or until the model fails to generate an in-vocabulary response in $k$ maximum attempts (in this case, $k=3$). We test several ablations varying prompt length, example types, and example selection. We provide results and other details from these ablations in the supplementary material (appendix). We repeat this sequence generation procedure on all $202$ event concepts, generating $2,276$ event sequences with an average length of $5.7$ labels per sequence.
The classical pattern mining algorithms (i.e. GSP and SPADE) we use to mine for new event patterns are both a priori-based approaches (Algorithm 1). Under both methods, given two sequences of the same event concept class
$\alpha=<a_{1},a_{2},..,a_{n}>\text{ and }\beta=<b_{1},b_{2},..,b_{n}>$
$\alpha$ is a subsequence of $\beta$ denoted as $\alpha\subseteq\beta$ iff there exists a set of values $1\leq j_{1}<j_{2}<...<j_{n}\leq m$ such that $a_{1}\subseteq b_{j1},a_{2}\subseteq b_{j2},...,a_{n}\subseteq b_{jn},$; then $\beta$ is a supersequence of $\alpha$. Given multiple such sequences and a support threshold, the task here is to find a set of frequent event subsequences.
2.4 Identifying Influencing Sets through Summary Markov Models
Learning about potential influencing events that lead to a given event type in a large set of event sequences is an important aspect of analyzing multivariate event sequences, i.e. events without time stamps, like the ones generated by our models. Classic $k_{th}$ order Markov chains capture these dynamics by modeling the probability of observing a particular event type given the preceding $k$ events in-sequence. The recent family of summary Markov models (SuMMs) Bhattacharjya et al. (2022) generalize other well known Markov models for event sequences by leveraging a function that summarizes historical event occurrences, and identifying the subset of event types that affect the probability of occurrence of event types of interest; this forms the influencing set.
We use the LLM-generated sequences to learn two types of SuMMs: binary SuMMs (BSuMMs) and ordinal SuMMs (OSuMMs). Briefly, in BSuMMs, it is only the presence or absence of an event in the relevant history that has an effect on the occurrence of other events, while in OSuMMs the order of the events is also taken into account. We refer the readers to Sections 3.3 and 3.4 in Bhattacharjya et al. (2022) for formal definitions and methods for learning SuMMs over event sequence collections.
Figure 1 shows examples of learned influencing sets using BSuMMs for refugee crisis and mass migration events. Two events mass migration and famine are identified as influencing events for refugee crisis. The model indicates that the occurrence of both mass migration and famine together has a $0.91$ probability of resulting in refugee crisis as a part of an event sequence. On the other hand the occurrence of mass migration in the absence of famine has a $0.31$ probability of resulting in refugee crisis. search strategy to efficiently discover the minimal influencing sets. We view the discovery of influencing sets from LLM-generated data as a mechanism of distilling complex symbolic knowledge from the output of LLMs.
3 Experimental Settings
We performed all our model inference related experiments on two NVIDIA V100 GPUs. We used the Huggingface library v4.26.14.26.14.26.14.26.1 Wolf et al. (2020) and publicly available model checkpoints.111huggingface.co/docs/transformers/model_doc/flan-t5 Classical pattern mining algorithms (GSP, SPADE) were implemented in Python through the use of spmf222pypi.org/project/spmf/ library v1.41.41.41.4.
For generation we use in-context learning under zero shot settings for all LLMs, with top-k sampling ($k=50$, in conjunction with top-p, with $p=0.95$).
To identify influencing events from Summary Markov Models (SuMMs), we use implementations from Bhattacharjya et al. (2022) and split the dataset (generated from LLMs) into train/dev/test sets ($70\%/15\%/15\%$) to generate results reported in Table 2. BSuMMs and OSuMMs were learned with hyperparameters of $\alpha=0.1$, $\gamma=0.5$ and a look-back ($\kappa$) window of $4$.
4 Evaluation
Open ended text generation in a task like event sequencing, with extremely sparse reference data, poses unique challenges to the evaluation of model outputs Wadhwa et al. (2023). The traditional scheme to evaluate discrete model outputs has been to calculate precision and recall for the generated outputs against a predefined reference test set. However, under open-world settings and particularly when a sparse KG like Wikidata is used as reference data, a missing causal relation between two event classes in a sequence may very well be a valid relation. Therefore, it is not possible to automatically derive an accurate measure of precision and recall purely using Wikidata as reference data.We take a multi-pronged evaluation approach to assess the quality and usefulness of the generated event sequences across multiple tasks.
4.1 Evaluation of Recall
Despite the sparsity of causal relations in Wikidata, one can still reasonably estimate recall through pairwise comparisons of generated event classes to the existing causal relations in Wikidata. We build a reference set of causal relations in Wikidata by curating a list of all the pairs of event concepts that have any of the several causal relations333www.wikidata.org/wiki/Wikidata:List_of_properties/causality in Wikidata in any direction, including has_cause (P828), has_immediate_cause (P1478), has_contributing_factor (P1479), and has_effect(P1542). Our results show that this is an effective mechanism of comparing recall across methods, as larger models that are expected to have better accuracy yield higher recall scores.
4.2 Evaluation of Precision
To overcome sparsity in reference data, prior work has generally relied on human evaluations Chiang and Lee (2023)for estimating precision, which entails looking at pairs of events and asking annotators whether or not the events in question have a causal relationship. Such a process for potentially thousands of event pairs (like in our case) can be very cost prohibitive. Furthermore, recent research Zhao et al. (2022); He et al. (2023) indicates that pretrained language models themselves might outperform lay human annotators such as those found on Amazon Mechanical Turk. For instance, He et al. (2023) demonstrated that labeling data under a few-shot chain-of-thought prompt (“explain-then-annotate” setting) surpasses crowdworker annotations on relevance assessments. Following these results, we use LLMs for evaluating precision and for selecting the most optimal model.
We propose evaluating precision for model selection as a binary classification task. Given two events $e_{1},e_{2}$, an evaluator model must evaluate whether $e_{1}$ reasonably leads to $e_{2}$ under a has_cause or has_effect relationship. To do this, we start with a large instruction-tuned model (in our case, Flan-T5-XXL ($11$B) Chung et al. (2022)) and adopt instructional in-context few shot prompting to classify whether or not $e_{1}$ reasonably leads to $e_{2}$, and for the model to provide a justification for its results. While we hypothesize that such an approach may yield noisy results, in a preliminary manual assessment this approach seems to yield reasonable outputs that are sufficient for evaluating correctness of model outputs (i.e. precision). We include a sample of model evaluated outputs in supplementary material, and select few examples in Appendix LABEL:sec_apx:examples.
Precision Evaluator Model (P) R F-1
Flan-T5-Large Flan-T5-XL Flan-T5-XXL
Flan-T5-Large ($783$M) Chung et al. (2022) 0.73 0.72 0.72 0.47 0.57
Flan-T5-XL (3B) 0.75 0.75 0.78 0.49 0.60
Open-Research LLaMA ($3$B) Touvron et al. (2023) 0.70 0.69 0.72 0.45 0.55
Flan-T5-XXL ($11$B) 0.79 0.79 0.81 0.54 0.65
[…]
4.3 Human Evaluation of Cause-Effect Prediction Accuracy of LLMs
In addition to this automated evaluation of our LLM-generated sequences, we carry out additional evaluation of our method on an existing popular knowledge graph – ATOMIC Hwang et al. (2021) and conduct human evaluation on a sample of data to support our approach. ATOMIC consists of event-centered pairs of instances of the form IsAfter (Y comes after X) and Causes (X causes Y). We leverage these instances as input prompts to the model and use human annotators to evaluate outputs. Specifically, we show human annotators anonymized model outputs and true references and elicit their preferences given a trigger event. We find that in an overwhelming majority of the cases, the models generate output that is semantically equivalent to the reference (even though there is no direct lexical alignment), or output that the humans prefer over the true reference. We provide example outputs of various evaluator models in the supplementary material (appendix).
4.4 Evaluation of Event Models
Apart from a qualitative manual analysis of the mined patterns, we follow Bhattacharjya et al. (2022) and use a metric that evaluates the quality of the learned model with respect to the data. Specifically, we use negative log loss as the evaluation metric, which is a popular metric for probabilistic prediction Bishop (2006). We posit that higher-quality event sequences result in better prediction ability of the models, and so better negative log loss scores for the models. Our results show that this is a valid assumption, as larger models that are expected to produce higher-quality sequences yield models with better prediction ability.
($\downarrow$) Data Generator Model BSuMMs OSuMMs LSTM
Flan-T5-Large (783M) Chung et al. (2022) -63.49 -84.29 -109.28
Flan-T5-XL (3B) -63.20 -92.65 -121.23
Open-Research LLaMA (3B) Touvron et al. (2023) -110.57 -101.29 -190.68
Flan-T5-XXL (11B) -57.99 -78.64 -108.89
Table 2: Negative log likelihood (lower magnitude is better) averaged over interest labels from LLM-generated (Flan-T5-XXL) event sequences. Lookback window ($k$) for LSTM was fixed to $5$. Best-performing data generator model scores are in bold.
5 Experimental Investigation and Results
We conducted a three-pronged quantitative and qualitative evaluation of our results. First, we evaluated LLM-generated event sequences in absolute terms through standard traditional metrics – precision, recall, and F1 score. Second, we gauge how probabilistic models like SuMMs fare when trained on LLM-generated data. Finally, we look at some qualitative examples of newly identified high support sequences (Appendix LABEL:sec_apx:examples) not already present in our base knowledge graph i.e. Wikidata.
5.1 Generated Event Sequences
We generate event sequences via LLMs through iterative in-context prompting (Figure 3). Table 1 summarizes our results from these experiments. We find that Flan-T5-XXL consistently outperforms its smaller counterparts, indicating that larger models generally fare better and benefit from their larger size (i.e. number of parameters). We posit this may also indicate that using larger models may continue to yield performance gains.
[…]
Input Instance: PersonX drops out of high school
Response 1: PersonX gets a job
Response 2: PersonX turns PersonX’s life around
Type: IsAfter
One of the responses above is the model (Flan-T5-XXL) output, while the other is the true reference. The human evaluators are then asked to answer the following questions–
  • Are both responses functionally similar?
  • Which response do you prefer?
  • Which response, if any, is completely irrelevant?
Based on the responses from human evaluators [^5] We observe strong inter-rater agreement with a Fleiss kappa, $\kappa=0.81$; conflicting response labels were aggregated through majority vote., we observe that humans found $65.82\%$ of response pairs functionally equivalent. That is, even though not lexically aligned, they meant the same thing. In $27.64\%$ of the instances the humans preferred the model-generated event instance over the true reference. In only $6.55\%$ of instances did the humans prefer the true reference over the model-generated output. [^6] Additional details on these experiments are available in the supplementary material (appendix). These results reinforce our underlying assertion that LLMs are capable of event-centered commonsense reasoning.
Precision Evaluator Models
We leverage our models for dual purposes – for generating and evaluating event sequences. Owing to this inherent conflict, we find it prudent to independently assess these evaluator models. To ensure robustness of our results, we use multiple evaluator models (Table 1) and find no significant difference between evaluated precision across different models used as evaluators.
5.2 Probabilistic Prediction via SuMMs and LSTM
Following the evaluation strategy implemented in Bhattacharjya et al. (2022), we focus on individual labels of interest and conduct an evaluation around probabilistic prediction. Consequently, we select negative log loss as the evaluation metric. Table 2 summarizes our results on the test set for both BSuMMs and OSuMMs, along with a simple LSTM baseline. We observe that models trained on event sequences generated from larger models (e.g. Flan-T5-XXL) fare better than the ones generated from their smaller counterparts (e.g. Flan-T5-Large). We treat this observation as a proxy for generated event sequence quality. That is, better quality sequences lead to better predictive models.
5.3 Qualitative Analysis
Figure 1 shows examples of mined patterns along with their support levels, as well as the probabilistic association between a number of influencing events from the output of BSuMMs. Appendix LABEL:sec_apx:examples provides two examples of newly mined sequences derived from pattern mining over LLM-generated sequences, and an additional two examples of influencing sets derived from the application of SuMMs. For each example, we show evidence supporting the accuracy of the extracted knowledge. We provide additional qualitative examples in supplementary material. Overall, these results show the high quality and usefulness of the distilled event sequence knowledge.
5.4 Error Analysis
We now describe two of the most common recurring errors in the LLM-generated outputs and SuMMs’ discovered influence sets.
No Sequences Generated
Out of the $202$ unique input event-triggers in our data, $8$ event-triggers yielded sequences of length $1$, i.e., no sequences. This means that the underlying model could not select a relevant event concept from the provided vocabulary that might follow the given trigger.
Illogical Influencing Sets
Due to the choice of the vocabulary of limited event concepts, we often find that the identified influencing sets (i.e., parent events) for some event concepts may not be correct or, at times, completely illogical. Qualitatively, we observe that this occurs with rare event types that do not appear often either in the underlying knowledge graph, or in the LLM-generated set of event sequences. Consequently, SuMMs fail to identify logical influencing sets. For example, for the event type United States Presidential Impeachment, BSuMM model identified unequal treaty as the influencing event.
6 Related Work
News Event Analysis
The primary applications of our work are around news event analysis and forecasting.  Zhao (2021) presents a taxonomy of different flavors of event prediction in the literature. Our target event prediction applications fall under the “Semantic Prediction” category, with time and location details not being of interest, and the primary goal being the prediction of “event profiles" such as event types. Seminal work in this area is the work of Radinsky et al. (2012) where causal relations between past events are extracted from text and then generalized to make predictions. Some recent work has explored the use of graph sequence mining Cekinel and Karagoz (2022) as a mechanism of extracting useful relations from large and noisy outputs of extraction. The application of building event sequence models over such outputs is largely unexplored. Our work explores an alternative approach of event sequence generation that is capable of generating longer, and potentially higher-quality sequences.
Event Sequence Extraction from Text
There is a wealth of literature on different methods of extracting sequences from textual corpora. Norambuena et al. (2023) present a comprehensive survey of automated methods of narrative extraction. Narratives contain several elements including events, participants (actors/protagonists), time, and space. The task of event detection itself is a highly challenging task and the topic of extensive research Chen et al. (2021); Xiang and Wang (2019), which has its root in the Topic Detection and Tracking (TDT) task Santana et al. (2023). This line of work started out as a DARPA-sponsored initiative with the same name Allan et al. (1998). Another closely related task is news timeline summarization Gholipour Ghalandari and Ifrim (2020). While pattern mining algorithms have been applied to the output of such extractions, e.g. for creation of “domain templates" Filatova et al. (2006), we are not aware of any attempts to use the extractions to construct event models for analysis or prediction.
Sequential Pattern Mining and Event Sequence Models
Sequential pattern mining and related approaches have been the subject of extensive research Mannila et al. (1997); Mabroukeh and Ezeife (2010); Mooney and Roddick (2013); Fournier-Viger et al. (2017). These algorithms take in a set of sequences (or “sequential records") with a set of unique events (or “items") and often a “minimum support" threshold, and return as output a ranked set of all frequent sequences in a given sequence collection (or database) meeting the minimum support threshold. There is also a long line of work on statistical and probabilistic models for analysis of different kinds of sequences. Our focus in this paper is on multivariate event sequences, i.e., sequences of various event types without timestamps. Markov models for sequences Raftery (1985); Begleiter et al. (2004) and long short-term memory (LSTM) Hochreiter and Schmidhuber (1997) models are examples of prediction models. We leverage a more recent family of models – Summary Markov models Bhattacharjya et al. (2022) – for some of the experiments in this paper that involve analyzing generated event sequences.
7 Conclusions
We evaluated the capabilities of large language models for the task of producing event sequences from concepts derived from a knowledge graph. We found that, through careful elicitation, LLMs are capable of producing high-quality event sequences that can be used for the downstream tasks of pattern mining and the construction of probabilistic event models such as SuMMs. We also proposed methods for effectively evaluating the outputs of LLMs with sparse reference sets and deriving useful signals for model selection through the use of existing traditional metrics. Our hope is that this work demonstrates the utility of large language models for extracting event-related information and helps researchers better navigate the complex interaction between event-related entities. While we have demonstrated some use-cases of the resulting dataset of event sequences (mining logical rules and patterns and extracting influencing event types), resulting datasets in this work and in other domains may themselves be a useful resource for researchers as one may draw upon them to inform or justify decision-making.